News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Sunday 8 August 1999 7.00am EST

The following is an overview of the events that occurred over the last 24 hours. Confirmed from additional state vector updates that the current Chandra orbit is its final operational orbit; Apogee is 139,188 km, perigee is 9,655 km, orbital period is 63 hours, 28 min, 43 sec. All values are well within specifications.

Successfully completed IPS deactivation.

Measured optical cavity pressure 8.1E-06 Torr. Pressure remains the same in the last 24 hours.

Propellant slosh tests confirmed that residual propellant slosh will not be a problem for PCAD.

Continued ACIS calibrations.

Further characterized the ISIM flexure thermal control anomaly to assess any possible affect on science observations scheduling.

Initiated Aspect Camera Assembly Full Field Search.

Events scheduled for the next 24 hours
  • Open ACIS door.
  • Begin ACIS background measurement calibrations.
  • Initiate installation of OBC-B On-orbit Load.
  • Begin ACA full field search evaluation and FID light characterization.
  • Begin Gyro Low Rate bias determination.

  • Return to Status Reports